# CP-Agent: Agentic Constraint Programming

### v2

@@ Abstract
Translating natural language into formal constraint models requires expertise in the problem domain and modeling frameworks. To investigate whether constraint modeling benefits from agentic workflows, we introduce CP-Agent, a Python coding agent using the ReAct framework with a persistent IPython kernel. Domain knowledge is provided through a project prompt of under 50 lines. The agent iteratively executes code, observes the solver’s feedback, and refines models based on the execution results.
We evaluate CP-Agent on CP-Bench’s 101 constraint programming problems. We clarified the benchmark to address systematic ambiguities in problem specifications and errors in ground-truth models. On the clarified benchmark, CP-Agent solves all 101 problems. Ablation studies indicate that minimal guidance outperforms detailed procedural scaffolding, and that explicit task management tools have mixed effects on focused modeling tasks.
@@ 1 Introduction
Constraint Programming (CP) is a declarative paradigm for solving combinatorial problems where users specify what constraints a solution must satisfy, rather than prescribing how to find it [26]. The primary limitation preventing wider adoption of CP remains the modeling bottleneck: translating natural language problem descriptions into formal constraint models requires deep expertise in both the problem domain and the modeling framework [5]. This challenge has been characterized as the “holy grail” of constraint programming [24].
@@
Large Language Models (LLMs) are now being applied to automate this translation process. The CP-Bench benchmark [14], introduced to evaluate LLM-driven constraint modeling capabilities, contains 101 diverse combinatorial problems sourced from various established CP community resources. These problems encompass optimization and satisfaction tasks, representing a variety of distinct constraint types. This benchmark addresses limitations of existing datasets, such as NL4Opt [17] and Logic Grid Puzzles [25], which lack the diversity and complexity necessary for evaluating constraint modeling capabilities.
@@
The CP-Bench study [14] evaluated multiple LLMs across three modeling frameworks. With direct prompting, the best models achieved up to 65% accuracy using Python-based frameworks (CPMpy, OR-Tools), while MiniZinc (a popular constraint modeling language) peaked around 50%. Applying inference-time methods (repeated sampling combined with self-verification) to CPMpy improved accuracy to approximately 70%.
@@
The choice between fixed and agentic workflows is well-established in the literature. As Wei et al. [29] demonstrated with chain-of-thought prompting, and subsequent work by Yao et al. [32] and Schick et al. [19] confirmed that one should use a fixed, single-shot workflow for well-scoped, short-horizon tasks, and prefer an agentic, tool-using loop as task complexity, multi-step reasoning, or external information needs increase. We hypothesize that constraint modeling from natural language falls into the latter category, requiring an agentic workflow with iterative refinement, hypothesis testing, and dynamic adaptation based on solver feedback. Our results confirm this hypothesis.
@@
To test this hypothesis, we introduce CP-Agent, a constraint modeling agent based on our general-purpose agentic-python-coder system. CP-Agent operates within a persistent IPython kernel [16], enabling stateful and iterative development. Rather than embedding domain-specific logic into its architecture, the agent is guided by a concise project prompt. This design allows the agent to execute code, observe runtime feedback from the constraint solver, and adjust models based on execution results.
@@
The underlying coding agent is also available as an MCP server [1]. The Model Context Protocol (MCP), now governed by the Linux Foundation, has become a de facto standard for connecting language models with external tools [8]. This integration enables interactive constraint modeling within MCP-compatible clients such as Claude Desktop or Cursor, where users can guide the agent through complex modeling decisions.
@@
On the CP-Bench dataset of 101 problems (with added clarifications for some ambiguous problems and output format requirements), CP-Agent solves all problems correctly. While our primary reference language model is Claude Sonnet 4.5, we also tested this approach with several other models from four different providers. Swapping only the project prompt enables the agent to adapt to new domains, as confirmed by follow-up work [23] on logic programming modeling tasks.
@@
We provide an experimental analysis of an explicit task management tool within our agentic coding workflow. The results are mixed: while task tracking helped on some previously failing problems, it also broke problems that worked without it. The net effect was negative (6 fewer problems solved) with significantly increased token usage, suggesting that for focused, single-file generation tasks, the overhead can distract as often as it organizes.
@@ 2 Related Work
CP-Agent aligns with the emerging neurosymbolic paradigm where neural language understanding is combined with symbolic constraint satisfaction [18]. The LLM-based agent translates natural language into formal constraints, while the CP solver provides a sound and complete search over the solution space.
Recent work has explored LLM agents for software engineering tasks. SWE-agent [31] resolves GitHub issues through a file-system-based Agent-Computer Interface with commands for navigating and editing repository files. CP-Agent focuses instead on single-script generation using a stateful IPython kernel, eliminating file I/O overhead for this class of problems. CodeAct [28] argues for using executable Python code as a unified action space for LLM agents, showing that this approach outperforms more restrictive formats. Our work applies this principle to constraint modeling, demonstrating that a general action space can solve specialized tasks by utilizing focused prompts. The iterative refinement in CP-Agent relates to automated program repair work. ChatRepair [30] introduced conversational repair, where an LLM iteratively refines patches based on test failures. RepairAgent [3] treats repair as an autonomous task with tools for code search, patching, and testing. Our agent performs similar iterative refinement but for generative tasks, translating natural language to constraint programs, rather than repairing existing code.
Benchmark quality affects evaluation validity. EvalPlus [13] demonstrated that popular benchmarks contain insufficient tests, with augmented test suites revealing previously undetected errors. ClassEval [4] addressed structural complexity through manual benchmark curation. Our work on CP-Bench addresses problem specification ambiguity, enabling more rigorous measurement of constraint modeling abilities.
@@ 3 Agentic Python Coder
CP-Agent is based on a generic Python coding agent equipped with a special prompt for constraint modeling. The coding agent implements a minimal yet capable coding agent using the ReAct framework [32]. The system architecture separates domain-agnostic agent capabilities from task-specific knowledge, enabling rapid adaptation to new problem domains through prompt engineering rather than code modification. The implementation of the generic coding agent is available from a GitHub repository and via PiPy for installation via “pip install agentic-python-coder” [20].
@@ 3.1 Architecture
Unlike software engineering agents that navigate complex file trees, this architecture targets single-script development, where all state is contained in a single file. The agent operates in a virtual execution space provided by the IPython kernel [16], which maintains a persistent state across multiple code executions. Variables, functions, and imports remain accessible throughout a session, eliminating file I/O overhead while enabling rapid hypothesis testing. The agent operates interactively via a Read–Eval–Print-Loop (REPL), and benefits from immediate feedback to all code manipulations. This design is well-suited for constraint programming models and similar focused generation tasks.
To manage this persistent state and guide the execution flow, we employ LangGraph [12]. This framework manages the agent’s workflow, enabling the development of stateful agents through durable execution and memory management. The framework supports the ReAct pattern [32] of interleaved reasoning and acting, allowing the agent to interact with the execution environment and adjust solutions through iterative cycles. The agent consists of three main components: the ReAct loop controller, which operates the reasoning-action cycle; a minimal tool suite focused on code generation and execution; and a project prompt that provides domain-specific guidance. The ReAct loop (Figure 1) manages the iterative cycle of reasoning about the task, selecting and executing tools, and observing results until the agent produces a final solution.
@@
Tool suite
The agent operates with a minimal, focused tool set. The core tool is python_exec(code: str), which executes Python code in the persistent kernel environment and returns stdout/stderr for solver feedback. Each execution builds on the previous state, allowing the agent to develop solutions incrementally. Calling save_code signals completion and writes the final solution to disk; a maximum iteration limit prevents runaway execution.
Kernel implementation
The system launches an IPython kernel as a separate subprocess, providing process isolation while enabling state persistence. All Python state resides in this long-running kernel, creating a stateful coding environment similar to an interactive session where each execution builds upon previous ones.
@@
Rather than having the agent execute pip install statements for adding required Python libraries or relying on pre-built container images, the system implements a custom kernel manager that wraps the standard IPython kernel launch. Using the uv package manager [2], the system injects requested packages directly into an ephemeral environment at kernel startup. This approach provides process-level isolation: each session runs in a clean environment with only the specified dependencies, while the kernel maintains state persistence across executions. The design decouples the agent’s reasoning from dependency mechanics, thereby eliminating the latency associated with runtime package installation.
Task management
The agent can use a task management tool to create and maintain an in-memory list of tasks, commonly referred to as a “todo list.” The tool call uses the format todo_write(todos: List[Dict[str, Any]]). Each task must contain id, content, status, and priority fields, with only one task marked as “in_progress” at any time.
@@
The agent connects to language models through OpenRouter, [^1] https://openrouter.ai providing access without direct API dependencies. This approach allows switching between different models by changing a single configuration parameter. The system streams responses with usage statistics and detailed tracking of tool calls. These logs provide insights into the agent’s problem-solving process and resource consumption.
@@
The agent operates with a three-layer prompt hierarchy. The system prompt (approximately 100 lines) defines general coding agent behavior and tool usage patterns without task-specific specialization, instructing the agent on how to reason about code, execute iterative development cycles, and utilize available tools effectively. The project prompt provides domain-specific knowledge (such as constraint modeling), examples, and best practices relevant to a particular problem class. The task prompt contains the specific problem to solve, typically read from a task.md file in the working directory or provided directly via a command line parameter. This separation enables the same agent infrastructure to tackle diverse problem domains by swapping project prompts while maintaining consistent operational behavior.
@@ 4 CP-Agent
For constraint programming, the generic coding agent is equipped with a domain-specific project prompt and uses CPMpy (see below) as the modeling framework. The approach relies on the agent’s ability to use general coding capabilities guided by domain expertise expressed in the project prompt, rather than specialized constraint programming logic embedded in the architecture. The agent can test simplified problem instances, verify constraint satisfaction on example data, implement custom validation logic, and refine models based on solver feedback. This flexibility, combined with the persistent execution environment, enables the agent to discover and correct modeling errors that would escape detection in single-pass generation approaches.
@@ 4.1 CPMpy
The agent uses CPMpy [7] as the constraint programming foundation. This Python-based library builds on NumPy for efficient handling of decision variables as arrays, allowing integration with the broader Python ecosystem. Its solver-agnostic design enables modeling problems once and solving them using various backends, including CP, MIP, SMT, and SAT solvers. The choice of Python matches findings from Michailidis et al. [14] that LLMs perform better with Python-based frameworks than domain-specific languages, likely due to the billions of lines of Python code in their training data.
@@ 4.2 Project Prompt
The project prompt (cpmpy.md) is under 50 lines and provides focused guidance for constraint modeling. The complete prompt is included in the GitHub repository [20].
The prompt contains five core rules: use CPMpy’s declarative modeling rather than implementing search algorithms, output only valid JSON using json.dumps(), always import the json module when outputting JSON, check the exact output format required, and manually test solutions to ensure they satisfy the problem requirements. A basic code template demonstrates the standard structure, and brief sections cover three essential CPMpy constraints (AllDifferent, sum, Circuit) and optimization via model.minimize() and model.maximize().
This minimal prompt resulted from an ablation study. Initial experiments utilized a substantially longer prompt, approximately 800 lines in length, which contained detailed modeling procedures, extensive problem-type catalogs, and debugging guidance. The shorter version achieved equal or better performance, suggesting that for modern LLMs with constraint programming knowledge in their training data, concise guidance focusing on common pitfalls suffices—excessive detail may limit problem-solving flexibility.
@@ 5 Benchmark Problems
The CP-Bench dataset [14] provides a benchmark for evaluating automated constraint modeling capabilities. It contains 101 constraint programming problems drawn from four established sources: CSPLib [6], CPMpy examples [7], models by Kjellerstrand [11], and problems from the APLAI course at KU Leuven. The problems span diverse combinatorial classes, including scheduling, assignment, sequencing, graph problems, and combinatorial design, with 30 optimization problems and 71 decision problems providing balanced coverage of constraint programming challenges.
Each problem is presented as a natural language description in Markdown format, along with input data and required output variables. Solutions must execute within a specified time limit and produce results in JSON format, using the specified variable names. CP-Bench includes ground-truth CPMpy models for all problems, enabling semantic validation through constraint satisfaction checking rather than relying solely on syntactic correctness or execution success. This semantic validation is essential for rigorous evaluation of automated modeling approaches. The ground-truth models are withheld from the agent and used solely for post-hoc verification of generated solutions.
@@ 5.1 Fixed Workflow Results
Michailidis et al. [14] evaluated fixed workflow approaches on CP-Bench, testing progressively detailed prompts (basic instructions, modeling guidelines, API documentation) and retrieval-augmented in-context learning. The highest accuracy of 70% was achieved by combining repeated sampling with self-verification using GPT-4-turbo. These fixed workflow approaches share a common limitation: they cannot test hypotheses, debug errors, or verify partial solutions.
@@ 5.2 Benchmark Corrections and Clarifications
Our initial evaluation attempts revealed issues with the original benchmark that prevented rigorous automated evaluation. The CP-Bench authors independently acknowledged222personal communication, August 2025 several problems, including ground-truth models with symmetry-breaking constraints that reject valid solutions, unclear output format requirements, and ambiguities in problem statements. Without clear specifications, evaluation conflates the ability to model constraint satisfaction problems with the ability to guess intended formats and interpret ambiguous descriptions. We added clarifications to isolate the actual modeling capability. These clarifications are necessary to maintain focus on constraint modeling rather than resolving ambiguity or making guesses.
Our benchmark updates fall into three categories. First, we clarified 31 ambiguous problem statements. For instance, Problem 080 (TSP) required specifying whether Euclidean distances should be rounded or truncated—different methods lead to different distance matrices and thus different optimal solutions. For Problem 013 (magic square), we corrected a misleading description that stated the goal is to find a “magic square” when the problem actually requires a “magic sequence.” Problem 060 (crossword) provided a visual grid and word list, but required explicit specification of slot lengths and intersection constraints to be solvable. Second, we added structured JSON schemas to all problems to clarify the required output format. Third, we corrected errors in 19 ground-truth models: removing symmetry-breaking constraints from 10 and fixing logic errors in 9 others.
The modifications do not make the constraint modeling problems easier—they remove confounding factors that obscure actual modeling capability. The underlying combinatorial structure and search space complexity remained unchanged; no changes hint at constraint formulations or modeling strategies. The success rates reported in the original CP-Bench paper were obtained using the original benchmark and ground-truth models. Direct comparison between our results and theirs would be inappropriate, as they measure different capabilities. With clear specifications, other methods would likely show higher success rates as well. We document all changes in supplementary materials and provide the clarified benchmark for fair comparison in future work.
6 Experiment

### v3

# Stefan Szeider Algorithms and Complexity Group TU Wien, Vienna, Austria

@@ Abstract
The translation of natural language to formal constraint models requires expertise in the problem domain and modeling frameworks. To explore the effectiveness of agentic workflows, we propose CP-Agent, a Python coding agent that uses the ReAct framework with a persistent IPython kernel. We provide the relevant domain knowledge as a project prompt of under 50 lines. The algorithm works by iteratively executing code, observing the solver’s feedback, and refining constraint models based on execution results.
We evaluate CP-Agent on 101 constraint programming problems from CP-Bench. We made minor changes to the benchmark to address systematic ambiguities in the problem specifications and errors in the ground-truth models. On the clarified benchmark, CP-Agent achieves perfect accuracy on all 101 problems. Our experiments show that minimal guidance outperforms detailed procedural scaffolding. Our experiments also show that explicit task management tools can have both positive and negative effects on focused modeling tasks.
@@ 1 Introduction
Constraint Programming is a declarative paradigm for solving combinatorial problems in which users specify the constraints a solution must satisfy, rather than prescribing how to find it [26]. The main obstacle preventing wider adoption of CP remains the modeling bottleneck. Converting natural language problem descriptions into formal constraint models requires deep expertise in both the problem domain and the modeling framework [5]. This challenge is referred to as the “holy grail” of constraint programming [24].
@@
In recent years, Large Language Models (LLMs) have been successfully applied to automate this translation process. We considered CP-Bench [14], a benchmark introduced to evaluate LLM-driven constraint modeling capabilities, which contains 101 diverse combinatorial problems sourced from various established CP community resources. These problems are used as benchmarks for optimization and satisfaction tasks, and include a variety of distinct constraint types. This benchmark addresses the limitations of existing datasets, such as NL4Opt [17] and Logic Grid Puzzles [25], which lack the diversity and complexity needed to evaluate constraint modeling capabilities.
@@
Michailidis et al. [14] evaluated several LLM-based workflows on CP-Bench. With direct prompting, Python-based frameworks (CPMpy, OR-Tools) achieve overall good accuracy up to 65%. Indeed, MiniZinc (a popular constraint modeling language) peaked at around 50%. Repeated sampling can be combined with self-verification to improve accuracy for Python-based frameworks to approximately 70%.
@@
The relationship between fixed and agentic workflows is now well understood. Chain-of-thought prompting [29] and subsequent work [32, 19] have been widely used in the literature. One distinguishes between fixed workflows that use well-scoped, short-horizon tasks and agentic workflows where the LLM runs in a loop and uses tools. We hypothesize that constraint modeling from natural language falls into the category of tasks that require an agentic workflow with iterative refinement, hypothesis testing, and dynamic adaptation based on solver feedback. Our experiments show that this is indeed the case.
@@
To test this hypothesis, we introduce CP-Agent, a constraint modeling agent that utilizes our general-purpose agentic-python-coder system. CP-Agent maintains a persistent IPython kernel [16], enabling stateful and iterative development. Instead of domain-specific logic embedded in its architecture, we use a concise project prompt as the guide. The agent can execute code, receive runtime feedback from the constraint solver, and adjust its constraint model accordingly.
@@
The underlying coding agent is available as an MCP server [1]. The Model Context Protocol (MCP), now governed by the Linux Foundation, is widely used for connecting language models with external tools [8]. With this integration, the CP agent can be used from any MCP-compatible client, including coding assistants like Cursor or AI chatbots like Claude Desktop.
@@
On the CP-Bench dataset of 101 problems (with added clarifications for some ambiguous problems and output format requirements), CP-Agent solves all problems correctly. We use Claude Sonnet 4.5 as the primary reference language model, though we also tested this with other models from different providers. Just by changing the prompt while keeping the agent architecture, we can adapt the system to different domains. This is exemplified by follow-up work [23] on modeling tasks for logic programming.
@@
In addition, we provide an empirical evaluation of using a task management tool within our agentic coding workflow. The results are mixed: task management is helpful for some but counterproductive for others; the net effect was adverse, with significantly increased token usage, suggesting that the overhead can distract as often as it organizes.
@@ 2 Related Work
Our approach follows the neurosymbolic paradigm, combining neural AI with symbolic constraint satisfaction [18]. The LLM-based agent converts natural language into formal constraints. The CP solver guarantees formal correctness of the computation.
LLM agents have been explored for software engineering tasks. SWE-agent [31] introduced the Agent-Computer Interface, a file-system-based approach for resolving GitHub issues that allowed navigating and editing repository files. Our approach is based on single-script generation using a stateful IPython kernel, which eliminates file I/O overhead for this class of problems. CodeAct [28] promotes executable Python code as a unified action space, and finds that this approach still outperforms more restrictive formats. We apply this general approach to constraint modeling, where a general action space can solve specialized tasks using focused prompts. The iterative refinement that we use in CP-Agent is related to automated program repair. ChatRepair [30] proposed conversational repair, where one iteratively refines patches based on test failures. RepairAgent [3] considers repair as an autonomous task with tools for code search, patching, and testing. Our approach is similar in spirit to iterative code repair, where refinement is used to correct errors. Still, we apply it to generative tasks, translating natural language into constraint programs rather than repairing existing code.
The validity of an evaluation depends on benchmark quality. EvalPlus [13] showed that popular benchmarks have insufficient tests, while augmented test suites reveal previously undetected errors. Only recently, ClassEval [4] proposed a new benchmark for structural complexity through manual curation. We address problem ambiguity in the original CP-Bench problems and propose a fixed version of the benchmark that allows more rigorous testing of constraint modeling capabilities.
@@ 3 Agentic Python Coder
Our CP-Agent is based on a generic Python coding agent, equipped with a special prompt for constraint modeling. Our coding agent implements the basic idea of ReAct [32] (Reason and Act) in a way that is both minimal and capable. The architecture can be adapted to new problem domains through prompt engineering rather than code modification. By this separation, domain-agnostic agent capabilities remain distinct from task-specific knowledge. Our implementation is publicly available on GitHub and via PyPI for installation via “pip install agentic-python-coder” [20].
@@ 3.1 Architecture
Unlike software engineering agents, which navigate complex file trees, we do not maintain multiple files; instead, development is focused on a single file. We process development, refinement, and execution in a virtual space provided by the IPython kernel [16], which runs as a subprocess. The kernel provides a persistent state across multiple code executions. This guarantees that variables, functions, and imports remain accessible throughout a session and enables rapid hypothesis testing without file I/O overhead. The agent proceeds in rounds, and each round consists of code manipulation followed by immediate feedback via the Read–Eval–Print Loop (REPL). This design is well-suited for developing constraint programming models and similar focused generation tasks.
The ReAct agent is implemented using LangGraph modules. LangGraph provides an important link between workflow management and durable execution, enabling the development of stateful agents. This even prevails when we consider the ReAct pattern [32] as interleaved reasoning and acting, which allows the agent to interact with the execution environment and adjust solutions through iterative cycles.
We will now explain the details of each component: the ReAct loop controller, which implements the reasoning-action cycle (using LangGraph [12]); three tools the agent can call; and a project prompt that provides domain-specific guidance. In each iteration of the ReAct loop, the agent reasons about the task, selects and executes tools, and observes results until a final solution is obtained and saved to a file (Figure 1).
@@
Tools
The main tool is python_exec(code: str), which runs Python code in the persistent kernel environment and returns stdout/stderr for solver feedback. Each call to this tool builds incrementally upon the previous calls. The function save_code transforms the solution into a file on disk, signaling completion. One could omit the save_code tool since file access is also possible via python_exec(code: str). However, having a dedicated tool for code saving has the advantage of easier file-path handling and makes it clear to the agent that calling this tool marks the end of its session. Optionally, we give the agent access to a task management tool to keep track of tasks in the form of a “to-do list.” The tool call has the format todo_write(todos: List[Dict[str, Any]]), where todos is the task list, and each entry specifies task attributes. By construction, each task must contain id, content, status, and priority fields, and only one task can be marked as “in_progress” at any time.
@@
We implement a custom kernel manager, which wraps the standard IPython kernel launch. Instead of having the agent run pip install commands or rely on prebuilt container images, we inject packages at startup. Using the uv package manager [2], we split package installation into an ephemeral environment that runs at kernel startup. The kernel resolves dependencies based on session configuration; each session runs in a clean environment with only the specified packages.
@@
The LLM that drives the agent is provided via the OpenRouter API. This allows us to use models from different model-makers through a uniform API and configuration, and with a single API key. We record all the agent’s steps and tool calls in log files. These logs provide valuable insight into the agent’s problem-solving process and resource consumption.
@@
The agent operates with the following three-fold prompt hierarchy: the system prompt (approximately 100 lines) specifies how the agent can be used to reason about code and execute iterative development cycles, defining general coding behavior and tool usage patterns without task-specific specialization. The project prompt combines domain-specific knowledge (such as constraint modeling) with examples and best practices relevant to this particular problem class. The concrete task to solve is in the task prompt, typically read from a task.md file in the working directory or provided directly via a command line parameter. The same agent infrastructure can tackle diverse problem domains by simply replacing the project prompt.
@@ 4 CP-Agent
For constraint programming, the generic coding agent is equipped with a domain-specific project prompt that instructs it to use CPMpy as the modeling framework. No specialized constraint programming logic is embedded in the architecture. The agent can create and test simplified problem instances and verify the constraint model on example data. The agent can even implement custom validation logic by refining models based on solver feedback. This flexible workflow, together with the persistent execution environment, provides an excellent environment for discovering and correcting modeling errors that would escape detection in single-pass generation approaches.
@@ 4.1 CPMpy
We use CPMpy [7] as the constraint programming foundation primarily due to its Python integration. This library is based on NumPy for efficient handling of decision variables as arrays, making it well-suited for integration with the broader Python ecosystem. Its solver-agnostic design allows a more refined selection of backends and hence enables modeling problems once and solving them using various solvers, including CP, MIP, SMT, and SAT. This design is in line with findings from Michailidis et al. [14] that LLMs perform better with Python-based frameworks than with domain-specific languages, likely due to the billions of lines of Python code in their training data.
@@ 4.2 Project Prompt
The project prompt (cpmpy.md), under 50 lines, provides succinct guidance for constraint modeling. The complete prompt is available online at the GitHub repository [20]. The prompt contains the following core rules: use CPMpy’s declarative modeling rather than implementing search algorithms, output only valid JSON using json.dumps(), always import the json Python module when outputting JSON, check the exact required output format, and manually test solutions to ensure they satisfy the problem requirements. A basic code template demonstrates the standard structure, and brief sections cover three essential CPMpy constraints (AllDifferent, sum, Circuit) and optimization via model.minimize() and model.maximize().
This minimal prompt is based on an ablation study. We initially experimented with a substantially longer prompt of approximately 800 lines, containing detailed modeling procedures, extensive problem-type catalogs, and debugging guidance. The shorter version performs on par with, or better than, the longer version across all performance metrics. This indicates that for modern LLMs with constraint-programming knowledge in their training data, concise guidance is sufficient; excessive detail may distract and consume input tokens without providing benefit.
@@ 5 Benchmark Problems
The CP-Bench collection [14] provides an excellent benchmark for testing automated constraint modeling capabilities. This collection comprises 101 constraint programming problems from four established sources: CSPLib [6], CPMpy examples [7], models by Kjellerstrand [11], and problems from the APLAI course at KU Leuven. The collection covers diverse combinatorial problem types, including scheduling, assignment, sequencing, graph problems, and combinatorial design. With 30 optimization and 71 decision problems, the collection provides a balanced coverage of constraint programming challenges.
Each problem in CP-Bench consists of a natural-language description in Markdown format, input data, required output variables, and a ground-truth model for verification. Solutions must output results in JSON format using the specified variable names. The ground-truth models enable semantic validation through constraint satisfaction checking rather than syntactic correctness or execution success. The ground-truth models are withheld from the agent and used solely for post-hoc validation.
@@ 5.1 Fixed Workflow Results
Michailidis et al. [14] evaluated fixed workflow approaches on CP-Bench, testing progressively detailed prompts including basic instructions, modeling guidelines, and API documentation, as well as retrieval-augmented in-context learning. Their results indicate that combining repeated sampling with self-verification using GPT-4-turbo achieves the highest accuracy of 70%. A general limit of these fixed workflow approaches is that they do not support testing hypotheses, debugging errors, or verifying partial solutions.
@@ 5.2 Benchmark Corrections and Clarifications
During testing, we encountered issues with some CP-Bench problems. The CP-Bench authors acknowledged these issues independently111personal communication, August 2025, including ground-truth models with symmetry-breaking constraints that reject valid solutions, unclear output format requirements, and ambiguities in problem statements. Without a clear problem specification, the modeling ability cannot be evaluated separately from format guessing. We therefore updated the CP-Bench problems to address these issues.
Our benchmark updates fall into three categories.
First, we clarified 31 ambiguous problem statements. For instance, Problem 080 (TSP) required specifying whether Euclidean distances should be rounded or truncated, which can lead to different distance matrices and thus different optimal solutions. For Problem 013 (magic square), we corrected a misleading description that stated the goal is to find a “magic square” when the problem actually requires a “magic sequence.” Problem 060 (crossword) provided a visual grid and word list, but required explicit specification of slot lengths and intersection constraints to be solvable.
Second, we added structured JSON schemas to all problems to clarify the required output format. We recommend this practice for any future extensions of the benchmark collection, as it eliminates trivial but annoying discrepancies, such as whether array indices are supposed to start with 0 or 1.
Third, we corrected errors in 19 ground-truth models: we removed symmetry-breaking constraints from 10 and fixed logic errors in 9 others.
While these updates do not make the problems easier, they remove confounding factors that obscure actual modeling capability. The underlying combinatorial structure and search space complexity remain the same; no changes hint at constraint formulations or modeling strategies. The success rates reported in the original CP-Bench paper were obtained without using our clarifications. A direct comparison between our results and theirs is not possible, as the fixed workflow can also benefit from the clarifications. The clarified benchmark is available online [22].
6 Experiments